4x2 = 35 - 4x
Thank you

Answer by AnlytcPhil(1810)

                      4x2 = 35 - 4x

Add 4x to, and subtract 35 from, both sides to get 0 on the right

            4x2 + 4x - 35 = 0 

Factor the left side:

         (2x + 7)(2x - 5) = 0

Use the 0-factor principle.  Set each factor on the left = 0.

Setting 2x + 7 = 0

            2x = -7

             x = -7/2

Setting 2x - 5 = 0

            2x = 5

             x = 5/2

There are two solutions, -7/2 and 5/2

Edwin
AnlytcPhil@aol.com